A ball is thrown horizontally from the top of a 80 m high cliff. If the initial horizontal velocity is 10.0 m/s, find

(i) how long it takes to reach the horizontal plane at the foot of the cliff.
1
Expert's answer
2020-07-02T17:15:16-0400

The key for solution is that the time we need to find does not depend on the horizontal velocity. It only depends on the height of the cliff because if there were no gravity, the rock would continue moving horizontally forever. Therefore, the acceleration due to gravity is the limiting factor:


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\cdot80}{9.8}}=4.05\text{ s}."